### v2

# Physical Commonsense Reasoning for Lower-Resourced Languages and Dialects: a Study on Basque

Abstract
Physical commonsense reasoning represents a fundamental capability of human intelligence, enabling individuals to understand their environment, predict future events, and navigate physical spaces. Recent years have witnessed growing interest in reasoning tasks within Natural Language Processing (NLP). However, no prior research has examined the performance of Large Language Models (LLMs) on non-question-answering (non-QA) physical commonsense reasoning tasks in low-resource languages such as Basque. Taking the Italian GITA as a starting point, this paper addresses this gap by presenting BasPhyCo, the first non-QA physical commonsense reasoning dataset for Basque, available in both standard and dialectal variants. We evaluate model performance across three hierarchical levels of commonsense understanding: (1) distinguishing between plausible and implausible narratives (accuracy), (2) identifying the conflicting element that renders a narrative implausible (consistency), and (3) determining the specific physical state that creates the implausibility (verifiability). These tasks were assessed using multiple multilingual LLMs as well as models pretrained specifically for Italian and Basque. Results indicate that, in terms of verifiability, LLMs exhibit limited physical commonsense capabilities in low-resource languages such as Basque, especially when processing dialectal variants.
Keywords: Physical Commonsense Reasoning, Multilingualism, Less-Resourced/Endangered Languages, Italian, Basque, dialects
[…]
Physical Commonsense Reasoning for Lower-Resourced Languages and Dialects: a Study on Basque
Jaione Bengoetxea, Itziar Gonzalez-Dios, Rodrigo Agerri
HiTZ Center - Ixa, University of the Basque Country UPV/EHU
{jaione.bengoetxea}@ehu.eus
Abstract content
1.  Introduction
Commonsense reasoning represents the human capacity to understand and manipulate real-world objects and their interactions. This domain has attracted considerable attention in Artificial Intelligence research in recent years Davis (2023); Sun et al. (2025). Physical commonsense reasoning, a specific subdomain, addresses events occurring in the physical world by capturing knowledge about everyday objects, their physical properties, and their potential uses and manipulations Bisk et al. (2020); Pensa et al. (2024a).
As a fundamental component of human intelligence, physical commonsense reasoning enables individuals to reason about their environment, anticipate future events, and navigate their surroundings. Recent research has increasingly examined the reasoning capabilities of LLMs, though such investigations have been conducted predominantly in English Bisk et al. (2020); Storks et al. (2021).
This paper focuses on Basque, specifically its Western dialect, as well as Italian, the source language of the dataset upon which our work is based on Pensa et al. (2024a, b). These low-resource languages provide valuable insight into LLM performance on complex physical-world reasoning tasks under data-limited conditions.
We manually translated the Italian dataset GITA into standard Basque and automatically adapted it into the Western dialect. The Western dialect was selected due to its peripheral status and documented linguistic distance from other Basque varieties, as established in dialectological research (Mitxelena, 1981). This linguistic divergence is corroborated by several NLP studies: Estarrona et al. (2023) identified Biscayan (Western) and Souletin as the most distinct among historical Basque dialects, while Bengoetxea et al. (2025) attributed the negative impact of the Western dialect on Natural Language Inference (NLI) performance to its distance from Standard Basque.
We evaluate model performance across three hierarchical reasoning tasks: (i) distinguishing plausible from implausible narratives (accuracy), (ii) identifying conflicting sentences within implausible narratives (consistency), and (iii) determining the physical states that render narratives implausible (verifiability). Our evaluation uses two multilingual LLMs alongside two Italian-pretrained models and one Basque-pretrained model, thereby examining current LLM knowledge of the physical world and human-object interactions.
To our knowledge, this represents the first investigation combining physical commonsense reasoning with Basque dialectal variation. Data and code are publicly available [^1] https://anonymous.4open.science/r/BasPhyCo-BBC9/README.md. Our investigation presents the following contributions:
  • The first publicly available non-QA physical commonsense reasoning dataset in Basque, including the first such dataset in a Basque dialect (Western).
  • The first evaluation of LLM performance on non-QA physical commonsense reasoning in a low-resource language such as Basque. Results indicate that, in terms of verifiability, LLMs exhibit limited physical commonsense capabilities in low-resource languages such as Basque, especially when considering dialectal variants.
  • A comprehensive evaluation of LLMs’ knowledge gaps when faced with physical commonsense reasoning for low-resource languages shows that this task is still challenging. Additionally, results with Basque language variation show that models pretrained for the target language seem to have a better ability to handle linguistic variation.
  • Fine-grained evaluation of physical states indicates that models have a general difficulty in correctly predicting these labels, Location, Edible, and Conscious states being particularly challenging.
2.  Related Work
Physical Commonsense
Recent research has tried to test physical commonsense knowledge of current LLMs. To this end, researchers have developed various datasets and benchmarks, including textual information Rajani et al. (2019); Bisk et al. (2020); Rajani et al. (2020); Storks et al. (2021); Aroca-Ouellette et al. (2021); Wang et al. (2023); Pensa et al. (2024a); Jeong et al. (2025), images Bakhtin et al. (2019); Hong et al. (2021); Liu et al. (2022); Meng et al. (2024), and videos Weihs et al. (2022); Yu et al. (2022); Motamed et al. (2025).
Datasets focusing on textual information have been generally presented as Question-Answering (QA) tasks, such as PIQA (Bisk et al., 2020). Some works have attempted to introduce other methodologies, TRIP (Storks et al., 2019), which is a physical commonsense reasoning benchmark composed of five-sentence stories. It evaluates models on three tasks: classifying stories as plausible or implausible, detecting the conflicting sentence, and identifying the physical states of objects involved.
The majority of the datasets in physical commonsense reasoning have been curated in English. Some exceptions include GITA (Pensa et al., 2024a) for Italian, a non-QA physical commonsense reasoning dataset based on TRIP, and EPiK Jeong et al. (2025) for Korean, which follows the PIQA dataset, while culturally adapting it to Korean.
To our knowledge, the sole existing resource for physical commonsense reasoning in Basque is a professionally translated version of the PIQA dataset (Baucells et al., 2025), which provides only the validation subset. Consequently, no Basque-language dataset for physical commonsense reasoning exists beyond the question-answering (QA) format.
Type Sentence 1 Sentence 2 Sentence 3 Sentence 4 Sentence 5
Plausible George filled the glass with water. George put the glass in the microwave. George turned on the microwave. The water got hot. George put a tea bag in the water.
Implausible (order) George put the glass in the microwave. George filled the glass with water. George turned on the microwave. The water got hot. George put a tea bag in the water.
Implausible (cloze) George filled the glass with water. George put the glass in the microwave. George turned on the microwave. The water got cold. George put a tea bag in the water
Table 1: Example of a story with its plausible and implausible versions.
Dialects and Reasoning
Regarding the use of dialects in commonsense reasoning, Lin et al. (2025) have very recently analyzed LLMs’ dialect robustness and fairness with Standardized English (SE) [^2] We use the terms the authors use in their papers. and African American Vernacular English (AAVE). They create the ReDial (Reasoning with Dialect Queries) dataset, a high-quality, end-to-end human-annotated SE-AAVE parallel dataset for reasoning tasks (algorithm, logic, math, and integrated reasoning) that contains over 1.2K parallel prompts in SE and in AAVE. An evaluation on LLM families (GPT, Claude, Llama, Mistral, Phi) shows lower performance when using dialectal prompts.
Pan et al. (2025) analyze dialectal bias on reasoning tasks through a multiple-choice question answering task, where they compare results in Standard American English (SAE)footnotemark: with results in 5 English dialects, such as Chicano, African American, or Indian English. The dataset was generated by applying grammatical perturbations to the original SAE multiple-choice benchmark using the Multi-VALUE package (Ziems et al., 2023). Results demonstrate that dialectal variation was the main reason for accuracy reductions of up to 20%.
Variation in Basque
Modern Basque dialects have been studied and categorized into a comprehensive representation of features by Zuazu (2008). In NLP, early works introduced a morpho-syntactically annotated corpus of Basque historical texts as an aid in the normalization process (Estarrona et al., 2020). Additionally, a corpus of syntactic variation of northern Basque dialects has been presented (Uria and Etxepare, 2012). More recently, Bengoetxea et al. (2025) presented the first manually created modern Basque dialect dataset for the evaluation of Natural Language Inference (NLI).
Finally, Basque dialects have also been considered in some dialectal benchmark works such as Alam et al. (2024) and Faisal et al. (2024), who presented benchmarks for MT with northern Basque dialects.
3.  Data
This study examines physical commonsense reasoning in Italian and Basque. We employed GITA (Pensa et al., 2024a), an Italian dataset derived from TRIP (Storks et al., 2019), and generated a Basque adaptation. GITA was selected as the foundation dataset due to its manual construction by a professional linguist with explicit attention to semantic coherence. Additionally, whereas TRIP incorporates compound sentences, GITA consists exclusively of simple sentences. This structural simplification reduces linguistic complexity and potential subjectivity, thereby isolating physical reasoning capabilities from confounding syntactic factors during model evaluation.
The following section introduces GITA and the process of its adaptation into both standard and dialectal Basque.
3.1.  GITA
GITA Pensa et al. (2024a) is an Italian physical commonsense evaluation dataset which consists of 356 5-sentence stories, out of which 117 are plausible, and 239 are implausible. The stories focus on concrete actions of the physical world, while mental actions such as “to think” or “to like” are avoided.
Two methods were used to create the implausible stories: (i) Order, where the order of two sentences was switched, and (ii) Cloze, where a plausible sentence has been substituted with an implausible one.
Consequently, individual sentences within the narratives are independently plausible but become implausible when placed with specific sentences in implausible narrative sequences. This design ensures that the reasoning task requires the use of the entire context.
In Table 1 we present an example translated into English. The plausible line contains the story with the logical sequence of events. In the implausible (order) example, the order of sentence 1 and 2 has been switched to make a non-logical and implausible story, and in the implausible (cloze) example, the adjective in sentence 3 has been changed from the original hot to cold, which makes no logical sense as the microwave heats water up.
3.2.  BasPhyCo
BasPhyCo is the first non-QA physical commonsense reasoning dataset for Basque, available in both standard and dialectal variants. BasPhyCo has been created by manually translating GITA from Italian to Standard Basque.
The translation process included a localization phase in which two linguists adapted cultural elements of GITA to align with Basque contexts. These adaptations included proper names and references to local meteorological agencies, among other culturally-specific elements. The translations adhered closely to standard Basque conventions, specifically excluding lexical items characteristic of Basque dialectal variants (guidelines will be provided in the Appendix upon acceptance).
$3.$  BasPhyCowest
The Standard Basque dataset was automatically converted to Western Basque using a few-shot prompting strategy implemented with the Latxa-3.1-Instruct model Sainz et al. (2025). Western Basque was selected for two reasons: (1) as a peripheral dialect, it exhibits substantial linguistic distance from Standard Basque, making it a valuable subject for comparative analysis; and (2) preliminary experiments with LLM-based automatic adaptation of GITA revealed a consistent tendency towards Western Basque generation. This methodology leveraged Latxa’s perceived tendency to generate Western dialect while accounting for its perceived divergence from standard Basque. The conversion prompt will be provided in the Appendix (upon acceptance).
Given that plausible and implausible story pairs contain identical sentences (with the exception of one sentence in cloze implausible narratives), the adaptation process grouped each plausible story with all corresponding implausible variants. The conversion prompt explicitly instructed consistent adaptation of repeated sentences across variants. This methodology ensured uniformity in the adapted narratives.
An example of this adaptation can be found in Table 2, where words like lorategia (garden) have been adapted to its Western form lorategixa, as well as auxiliary verbs such as du has been adapted to dau.
Standard Dialectal
Jonek lorategi handi bat dauka. Elur lorategian dago. Jonek lorategiko atea ireki du. Elurrek alde egin du. Lorategia hutsik dago. Jonek lorategi handi bat dauko. Elur lorategixen dago. Jonek lorategiko atie zabaldu dau. Elurrek ospa egin dau. Lorategixe hutsik dago.
Table 2: Example of a story adapted from Standard Basque to Western Basque.
A native professional linguist validated the automatic adaptations to assess overall quality (including minor formatting issues mitigated through prompt engineering) and identify dialectal adaptation errors. The subsequent subsections detail the findings from this manual inspection.
3.3.1.  Correct Adaptations
During the manual evaluation step, different types of dialectal linguistic modifications were identified.
Lexical features
Some lexical changes found to correspond to the Western dialect include itzali > amatatu (to switch off), galtzak > prakak (trousers) or jolas egin > olgetan egin (to play), to name a few. Not only that, but many words have also displayed Western phonology features, such as ordulariXE > ordulariA (clock) or salda > saldea (soup).
Morphosyntactic features
Some common Western morphosyntactic characteristics include the comitative (norekin, with what/who) case marker, which in Standard Basque is marked with -KIN, while in the Western dialect this case is represented with the termination -GAZ, as in the following example: aterkiareKIN > aterkixeGAZ (with the umbrella).
In terms of auxiliary verb forms, the majority of them have been adapted into the Western dialect, such as da > dau, nuen > neban, ditut > dodaz, to name but a few.
3.3.2.  Incorrect Adaptations
During the manual inspection of the adapted dialectal sentences, we found the following errors.
Lexical deviations
Some sentences contained made-up words that looked like dialectal words, such as mugikorra (phone, standard) > *mobillora or tomate (tomato, standard) > *totame. These lexical adaptations are not part of the Western dialectal vocabulary and could be considered examples of model hallucination. However, they represent a very minimal part of the whole dataset.
Additionally, some words contained changes that mimic Western dialectal phonology (e.g. baten > *paten, sagar > *saga), but are not in fact a part of Western dialectal phonological changes.
Morphosyntactic deviations
Although sentences generally follow dialectal morphosyntactic patterns, some outputs are not aligned with known dialectal features. For instance, some sentences with missing or additional ergative markers were found: the sentence *TeknikarixaK ez dau oraindiño etorri [^3] Translation: the technician has not arrived yet., has an extra ergative marker -K, as intransitive verbs do not need this marker.
Additionally, some sentences contained verb concordance mismatches, such as *indiolarrak hartu dau [^4] Translation: [someone] has taken the turkey., where the noun the verb is referring to is plural, but the verb form is singular. Thus, the preferred form would be indiolarrak hartu dauz
The observed morphosyntactic divergences are not attested in dialectal corpora, suggesting they stem from the model’s generalization errors rather than dialectal norms.
Figure 1: The number of unique words in BasPhyCo (left) and BasPhyCowest (right), as well as the overlap of both datasets (middle). Additionally, some examples from each dataset.
In Figure 1, we illustrate differences and similarities between the Standard and Western Basque datasets, highlighting both their lexical overlaps and divergences. While a portion of the vocabulary is shared between the two varieties, the analysis reveals that there is a substantial part of the lexicon that differs. We additionally present a series of contrastive examples that exemplify the most salient lexical and orthographic variations across the datasets.
4.  Experimental Setup
This section presents the three evaluated tasks and their associated metrics, followed by a description of the selected models and evaluation framework.
4.1.  Task description
Our setup is based on GITA4CALAMITA, a GITA version which was adapted to work with generative LLMs for the CALAMITA shared task (Pensa et al., 2024b). This approach evaluated three different tasks, which were based on the mirroring of the human reasoning, from the shallowest to the deepest. The evaluated tasks are the following:
  • Story classification determines if the story is plausible or not. Continuing with the example in Table 1, the plausible story should be classified as plausible and the other two as implausible.
  • Conflict detection involves identifying sentence pairs where the story becomes implausible. The conflicting sentences in the example of implausible-order in Table 1 are sentence 1 and 2, since once George puts the the glass in the microwave, it is not logical to fill it with water.
  • Physical state classification recognizes the involved physical states in the conflicting sentences of implausible stories. In the case of the example implausible-cloze in Table 1, the involved physical state is the temperature.
As in GITA4CALAMITA, we restrict the physical states to 14: location, conscious, dressed, wet, exist, clean, power, functional, in pieces, open, temperature, solid, occupied, and edible.
4.1.1.  Data Annotation
We adopt the annotation from Pensa et al. (2024b), which was manually revised by a professional linguist. Some minor annotation errors were detected and corrected, such as occasional mislabeling between cloze and order story types.
The following is an example from the dataset and its annotation. Some relevant fields include Type, which can be Null for plausible stories, and Order or Cloze for implausible ones; Confl_sents and Confl_pairs, that indicate which sentences make the story implausible.
"0-C0": {
"story_id": 0,
"type": "cloze",
"sentences": [
"Mikelek hozkailua ireki du.",
"Mikelek esnea hartu du.",
"Mikelek katilua hartu du.",
"Mikelek goilara hartu du.",
"Mikelek goilara katiluan sartu du."
"length": 5,
"example_id": "0-C0",
"plausible": false,
"breakpoint": 1,
"confl_sents": [0],
"confl_pairs": [0, 1]
Test data Model Accuracy Consistency Verifiability
Italian Llama-3.1-8B-It 72.47 29.29 14.23
Llama-3.1-70B-It 87.64 65.27 36.40
Gemma-2-9B-It 71.91 35.98 16.74
Gemma-2-27B-It 67.42 30.96 15.06
Latxa-3.1-8B-It 67.42 23.85 13.39
Latxa-3.1-70B-It 85.96 60.25 40.17
Minerva-7B-It 38.20 1.67 0.00
LlaMAntino-3-8B-It 58.99 18.41 7.95
Basque (Standard) Llama-3.1-8B-It 57.30 11.30 5.02
Llama-3.1-70B-It 84.83 47.70 26.78
Gemma-2-9B-It 66.01 25.10 7.53
Gemma-2-27B-It 62.36 24.27 8.37
Latxa-3.1-8B-It 65.45 23.43 8.79
Latxa-3.1-70B-It 81.46 48.12 30.54
Minerva-7B-It 36.52 2.09 0.42
LlaMAntino-3-8B-It 37.64 3.77 1.67
Basque (Western) Llama-3.1-8B-It 51.12 9.62 4.18
Llama-3.1-70B-It 74.16 35.56 17.57
Gemma-2-9B-It 64.61 21.34 5.44
Gemma-2-27B-It 57.87 18.41 6.69
Latxa-3.1-8B-It 63.48 17.99 9.21
Latxa-3.1-70B-It 80.34 46.86 28.03
Minerva-7B-It 38.48 2.51 0.42
LlaMAntino-3-8B-It 36.24 2.93 1.67
Table 3: Overall results for Story Classification, Conflict Detection and Physical State Classification, measured by accuracy, consistency and verifiability, respectively.
4.2.  Metrics
To evaluate model performance, we adopt a tiered evaluation framework (Storks et al., 2021; Pensa et al., 2024b). In this setup, each task is evaluated conditionally on the success of the previous one, forming a crescendo of increasingly demanding reasoning requirements. Specifically, only the correctly solved instances from one level are used as input to the next. Accordingly, we adopt three complementary metrics for the three evaluated tasks:
  • Accuracy: Quantifies the proportion of the correctly identified plausible and implausible stories. This metric will be used in the story classification task.
  • Consistency: Measures the proportion of the correctly identified plausible sentences and the conflicting sentence in the implausible stories. The aim of this measure is to check models’ consistency when recognizing conflicts. Thus, this metric will be used to evaluate the conflict detection task.
  • Verifiability: Evaluates the proportion of the correctly identified plausible sentences, the conflicting sentence and underlying physical states. This shows that the detected conflict can be validated because the underlying implausible change of physical states has been correctly understood. This last metric will be used to evaluate the physical state classification task.
4.3.  Evaluation Setting
We have evaluated our task on generative models, as previous works that evaluated discriminative models (Storks et al., 2019; Pensa et al., 2024a) were outperformed by generative models (Pensa et al., 2024b). The evaluation for the three tasks is implemented on EleutherAI’s Language Model Evaluation Harness framework v0.4.9 (Gao et al., 2024). This system enables the evaluation of generative LLMs and tasks in a reproducible, automated, and systematic way. The experiments were carried out in the few-shot setting specified by Harness.
We evaluated all tasks across the three test datasets representing Italian and Standard and Western Basque (Section 3). The evaluation employed four multilingual models, Llama-3.1 of 8B and 70B parameters Dubey et al. (2024) and Gemma-2 9B and 27B parameters Team et al. (2024), alongside language-specific models pretrained on Italian (Minerva-7B Orlando et al. (2024) and LlaMAntino-3-8B Polignano et al. (2024)) and Basque, namely, Latxa-3.1-8B and Latxa-3.1-70B Sainz et al. (2025). All models were instruction-tuned variants.
Language Model Accuracy Consistency Verifiability
Order Cloze Plausible Order Cloze Order Cloze
Italian Llama-3.1-8B-It 64.75 81.20 71.79 14.75 44.44 4.92 23.93
Llama-3.1-70B-It 88.52 95.73 78.63 55.74 74.36 25.41 47.01
Gemma-2-9B-It 49.18 81.20 86.32 15.57 57.26 2.46 31.62
Gemma-2-27B-It 34.43 78.63 90.60 9.02 53.85 4.10 26.50
Latxa-3.1-8B-It 55.74 68.38 78.63 13.11 35.04 6.56 20.51
Latxa-3.1-70B-It 84.43 94.87 78.63 45.90 75.21 25.41 55.56
Minerva-7B-It 13.11 13.68 88.89 0.00 3.42 0.00 0.00
LlaMAntino-3-8B-It 38.52 61.54 77.78 7.38 29.91 1.64 14.53
Basque (Standard) Llama-3.1-8B-It 46.72 50.43 75.21 8.20 14.53 2.46 7.69
Llama-3.1-70B-It 79.51 88.03 87.18 31.97 64.10 18.03 35.90
Gemma-2-9B-It 51.64 72.65 74.36 13.11 37.61 3.28 11.97
Gemma-2-27B-It 40.16 70.94 76.92 9.84 39.32 2.46 14.53
Latxa-3.1-8B-It 48.36 67.52 81.20 10.66 36.75 4.10 13.68
Latxa-3.1-70B-It 76.23 87.18 81.20 36.07 60.68 18.03 43.59
Minerva-7B-It 11.48 16.24 82.91 0.82 3.42 0.00 0.85
LlaMAntino-3-8B-It 6.56 12.82 94.87 0.00 7.69 0.00 3.42
Basque (Western) Llama-3.1-8B-It 34.43 44.44 75.21 4.10 15.38 1.64 6.84
Llama-3.1-70B-It 67.21 76.92 78.63 20.49 51.28 6.56 29.06
Gemma-2-9B-It 66.39 68.38 58.97 13.93 29.06 3.28 7.69
Gemma-2-27B-It 42.62 59.83 71.79 6.56 30.77 1.64 11.97
Latxa-3.1-8B-It 50.00 64.96 76.07 8.20 28.21 3.28 15.38
Latxa-3.1-70B-It 78.69 83.76 78.63 32.79 61.54 17.21 39.32
Minerva-7B-It 22.13 17.95 76.07 1.64 3.42 0.82 0.00
LlaMAntino-3-8B-It 4.92 8.55 96.58 0.00 5.98 0.00 3.42
Table 4: Fine-grained examples for all three metrics.
5.  Results
We present the results for the three tasks in Table 3, for Italian (GITA), Standard Basque (BasPhyCo) and Western Basque (BasPhyCowest).
Italian
The multilingual Llama-3.1-70B-It model achieved the highest performance in accuracy and consistency metrics, while Latxa-3.1-70B-It outperforms other models in terms of verifiability. Conversely, Italian-pretrained models (Minerva-7B-It and LlaMAntino-3-8B-It) yield the lowest performance across all evaluated tasks, with Minerva-7B-It showing notably inferior results compared to LlaMAntino-3-8B-It.
Notably, Basque-trained Latxa models outperformed Italian-specific models when evaluated on Italian data. Specifically, the smaller Latxa-8B-It model, despite being comparable in size to the Italian models, consistently surpassed LlaMAntino-3-8B-It across all tasks. This performance advantage can be attributed to Latxa’s continual pretraining approach Etxaniz et al. (2024), which effectively mitigates catastrophic forgetting from its base model, Llama-2.
Standard Basque
While Llama-3.1-70B-It obtains the highest accuracy score for story classification (84.83 vs 81.46), the Basque pretrained model Latxa-3.1-70B-It has higher scores for the other two more fine-grained metrics, consistency (47.70 vs 48.12) and verifiability (26.78 vs 30.54) respectively.
Western Basque
Latxa-3.1-70B-It obtains the highest results across all metrics. Llama’s performance drop from standard to dialectal data is worth mentioning, as all three metrics undergo important drops (84.83 vs 74.16 for accuracy, 47.70 vs 35.56 for consistency, 26.78 vs 17.57 for verifiability). With Latxa, although there is a performance drop from standard to dialectal, the drop is not nearly as dramatic (81.46 vs 80.34, 48.12 vs 46.86, 30.54 vs 28.03). These results highlight the importance of pretraining in the target language, as it appears to facilitate more fine-grained linguistic competence and enhance robustness to language variation.
Overall
LLMs demonstrate notably poor performance in predicting verifiable instances for low-resource languages, with performance degrading further when applied to dialectal data. Regarding task-specific performance, Llama-3.1-70B-It exhibited optimal results in story classification for Italian and Standard Basque, whereas Latxa-3.1-70B-It demonstrated superior consistency and verifiability, particularly for Standard and Western Basque. These results indicate that pretraining on target language data yields more substantial improvements in complex reasoning tasks. Additionally, Latxa-3.1-70B-It achieved the highest performance in verifiability, which is the most cognitively demanding reasoning task across all evaluated languages.
Finally, the drop from the shallowest to the deepest reasoning task for all models is to be highlighted. Table 3 shows substantial performance degradation, especially in the physical state classification task (verifiability). These findings indicate that, although some models are able to identify implausible stories, providing explanations for their implausibility presents a considerably more challenging task. This will be further discussed in Section 6.
6.  Discussion
In this section, we focus on more fine-grained results, as the three metrics have been specifically computed for the different types of implausible stories (order and cloze). The aim of this analysis is to identify any possible biases that the models could have towards implausible story types.
The results for all three metrics, as well as for the different types of implausible stories, are presented in Table 4. The main finding indicates that order implausible stories consistently yield lower scores than cloze implausible stories across all metrics, models, and languages. This pattern suggests that the models exhibit stronger reasoning capabilities when confronted with a conflicting sentence within a narrative sequence, compared to cases where implausibility arises solely from the reordering of sentences. These results are consistent with the findings reported by Pensa et al. (2024b).
P. state total Italian Basquestd Basquewest Avrg
Open 88 20.45 19.32 12.50 17.42
Functional 53 26.41 15.09 16.98 19.49
Exist 47 27.66 23.40 21.28 24.11
Power 36 41.67 36.11 13.63 30.47
In pieces 35 42.86 25.71 31.43 33.33
Location 33 15.15 6.06 6.06 9.09
Edible 22 4.54 0.00 4.54 3.03
Conscious 13 7.69 7.69 15.38 10.25
Temperature 12 50.00 50.00 41.67 47.22
Wet 7 42.86 58.57 14.28 38.57
Solid 5 80.00 60.00 40.00 60.00
Wearing 3 0.00 0.00 0.00 0.00
Clean 1 0.00 0.00 0.00 0.00
Occupied 1 100.00 100.00 100.00 100.00
Table 5: Verifiability results per physical state. These results are for Latxa-3.1-70B-It, the model with the highest verifiability results for Italian, Standard and Western Basque.
Italian and Standard Basque seem to follow similar patterns. Llama-3.1-70B-It obtained the highest results in the majority of the tasks and story types, only being surpassed by Latxa-3.1-70B-It in consistency and verifiability cloze story types. This suggests that, for Italian and Standard Basque, while Llama obtains higher results in shallower reasoning tasks (story classification), Latxa seems to perform slightly better in reasoning tasks involved physical state classification (verifiability).
Regarding the results for Western Basque, Latxa-3.1-70B-It outperforms all other models, including both multilingual models and those pretrained for Italian, following general results in Table 3.
Furthermore, the general decrease in performance observed for Llama-3.1-70B-It compared to the standard Basque results highlights the need for multilingual language models that could better handle Basque dialectal variation.
Finally, Latxa-3.1-70B-It consistently obtains high verifiability results for both order and cloze types, which is the metric that measures how much physical states are predicted correctly. This suggests Latxa’s capacity to deal with deeper reasoning tasks such as physical state classification.
Per Physical State Label Verifiability
In Table 5, we report the verifiability results for each physical state label across Italian, as well as Standard and Western Basque. Labels represented by fewer than ten instances are excluded from the following analysis, due to potential sampling bias. Consequently, the subsequent analysis focuses exclusively on those physical states with sufficient representation (i.e., more than ten instances), ensuring more reliable and interpretable results.
Overall, the findings indicate that no particular physical state is consistently easier to predict than the others. In general, performance across categories remains relatively low, highlighting both the intrinsic complexity of this reasoning task and the current limitations of LLMs in capturing nuanced physical state distinctions.
The predictions of Location, Edible, and Conscious states appear to be particularly challenging, as reflected by their comparatively lower verifiability scores. These results suggest that such categories may involve subtleties that LLMs struggle to capture effectively, possibly due to their dependence on implicit world knowledge.
7.  Conclusion
This paper introduces a novel dataset for evaluating physical commonsense reasoning in Basque and its Western dialect. The dataset was derived from GITA, a manually curated Italian corpus, which underwent manual translation and localization into Standard Basque. Subsequently, the Standard Basque data were automatically adapted to the Western Basque dialect, followed by manual post-editing to ensure accuracy and minimize errors.
We have carried out a suite of experiments to see how multilingual and language-specific LLM perform on the tasks of physical commonsense reasoning. To our knowledge, this is the first evaluation of non-QA physical commonsense reasoning in low-resourced languages such as Basque and its dialectal varieties. To that end, we have followed a tiered strategy with three tasks of different depth levels: story classification, conflict detection and physical state classification. The results show the LLMs ability to predict verifiable instances is generally low, which highlights the need for further research in the field of physical commonsense reasoning. Further analysis has indicated that identifying implausible instances is more complex when the only change is sentence order. Finally, physical state classification remains a particularly challenging task.
This work establishes a baseline evaluation framework for commonsense reasoning in low-resource languages and dialectal varieties. Future research directions include extending the dataset to additional languages and dialects.
Limitations
The physical commonsense reasoning dataset that we present in this work can be culturally localized, reflecting the norms and logic of certain communities, and may need to be adapted to other cultures in order to be applicable in other contexts.
Additionally, the size of our dataset is currently limited. Expanding this test data as well as building a training set could alleviate this issue.
Finally, it is important to recognize the inherent bias of Basque LLMs toward Western Basque. Current models show a strong tendency to generate Western Basque features, indicating that their training data and modeling are heavily aligned with this dialect. Expanding this ability to other dialects could enable the analysis of additional variations.
Acknowledgments
This work has been supported by the HiTZ center and the Basque Government (Research group funding IT-1805-22). Jaione Bengoetxea is funded by the Basque Government pre-doctoral grant (PRE_2024_1_0028). We also acknowledge the following MCIN/AEI/10.13039/501100011033 project: DeepMinor (CNS2023-144375) and European Union NextGenerationEU/PRTR.
Bibliographical References
  • M. M. I. Alam, S. Ahmadi, and A. Anastasopoulos (2024) CODET: A Benchmark for Contrastive Dialectal Evaluation of Machine Translation. In Findings of the Association for Computational Linguistics: EACL 2024, Y. Graham and M. Purver (Eds.), St. Julian’s, Malta, pp. 1790–1859. External Links: Link Cited by: §2.
  • S. Aroca-Ouellette, C. Paik, A. Roncone, and K. Kann (2021) PROST: Physical Reasoning about Objects through Space and Time. In Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021, C. Zong, F. Xia, W. Li, and R. Navigli (Eds.), Online, pp. 4597–4608. Cited by: §2.
  • A. Bakhtin, L. van der Maaten, J. Johnson, L. Gustafson, and R. Girshick (2019) PHYRE: A New Benchmark for Physical Reasoning. Advances in Neural Information Processing Systems 32. Cited by: §2.
  • I. Baucells, J. Aula-Blasco, I. de-Dios-Flores, S. Paniagua Suárez, N. Perez, A. Salles, S. Sotelo Docio, J. Falcão, J. J. Saiz, R. Sepulveda Torres, J. Barnes, P. Gamallo, A. Gonzalez-Agirre, G. Rigau, and M. Villegas (2025) IberoBench: A Benchmark for LLM Evaluation in Iberian Languages. In Proceedings of the 31st International Conference on Computational Linguistics, O. Rambow, L. Wanner, M. Apidianaki, H. Al-Khalifa, B. D. Eugenio, and S. Schockaert (Eds.), Abu Dhabi, UAE, pp. 10491–10519. External Links: Link Cited by: §2.
  • J. Bengoetxea, I. Gonzalez-Dios, and R. Agerri (2025) Lost in Variation? Evaluating NLI Performance in Basque and Spanish Geographical Variants. In Proceedings of the 29th Conference on Computational Natural Language Learning, G. Boleda and M. Roth (Eds.), Vienna, Austria, pp. 452–468. External Links: Link, Document, ISBN 979-8-89176-271-8 Cited by: §1, §2.
  • Y. Bisk, R. Zellers, J. Gao, Y. Choi, et al. (2020) PIQA: Reasoning about Physical Commonsense in Natural Language. In Proceedings of the AAAI conference on artificial intelligence, Vol. 34, pp. 7432–7439. Cited by: §1, §1, §2, §2.
  • E. Davis (2023) Benchmarks for Automated Commonsense Reasoning: A Survey. ACM Comput. Surv. 56 (4). External Links: ISSN 0360-0300 Cited by: §1.
  • A. Dubey, A. Jauhri, A. Pandey, A. Kadian, A. Al-Dahle, A. Letman, A. Mathur, A. Schelten, A. Yang, A. Fan, et al. (2024) The llama 3 herd of models. arXiv preprint arXiv:2407.21783. Cited by: §4.3.
  • A. Estarrona, I. Etxeberria, R. Etxepare, M. Padilla-Moyano, and A. Soraluze (2020) Dealing with Dialectal Variation in the Construction of the Basque Historical Corpus. In Proceedings of the 7th Workshop on NLP for Similar Languages, Varieties and Dialects, M. Zampieri, P. Nakov, N. Ljubešić, J. Tiedemann, and Y. Scherrer (Eds.), Barcelona, Spain (Online), pp. 79–89. External Links: Link Cited by: §2.
  • A. Estarrona, I. Etxeberria, M. Padilla-Moyano, and A. Soraluze (2023) Measuring Language Distance for Historical Texts in Basque. Procesamiento del Lenguaje Natural 70, pp. 53–61. Cited by: §1.
  • J. Etxaniz, O. Sainz, N. Perez, I. Aldabe, G. Rigau, E. Agirre, A. Ormazabal, M. Artetxe, and A. Soroa (2024) Latxa: an open language model and evaluation suite for Basque. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), L. Ku, A. Martins, and V. Srikumar (Eds.), pp. 14952–14972. Cited by: §5.
  • F. Faisal, O. Ahia, A. Srivastava, K. Ahuja, D. Chiang, Y. Tsvetkov, and A. Anastasopoulos (2024) DIALECTBENCH: A NLP Benchmark for Dialects, Varieties, and Closely-Related Languages. ArXiv abs/2403.11009. External Links: Link Cited by: §2.
  • L. Gao, J. Tow, B. Abbasi, S. Biderman, S. Black, A. DiPofi, C. Foster, L. Golding, J. Hsu, A. Le Noac’h, H. Li, K. McDonell, N. Muennighoff, C. Ociepa, J. Phang, L. Reynolds, H. Schoelkopf, A. Skowron, L. Sutawika, E. Tang, A. Thite, B. Wang, K. Wang, and A. Zou (2024) The Language Model Evaluation Harness. Zenodo. External Links: Document, Link Cited by: §4.3.
  • Y. Hong, L. Yi, J. Tenenbaum, A. Torralba, and C. Gan (2021) PTR: A Benchmark for Part-based Conceptual, Relational, and Physical Reasoning. Advances in Neural Information Processing Systems 34, pp. 17427–17440. Cited by: §2.
  • J. Jeong, D. Lee, D. Lee, and H. Yu (2025) Everyday Physics in Korean Contexts: A Culturally Grounded Physical Reasoning Benchmark. arXiv preprint arXiv:2509.17807. Cited by: §2, §2.
  • F. Lin, S. Mao, E. La Malfa, V. Hofmann, A. de Wynter, X. Wang, S. Chen, M. J. Wooldridge, J. Pierrehumbert, and F. Wei (2025) Assessing Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pp. 6317–6342. Cited by: §2.
  • X. Liu, D. Yin, Y. Feng, and D. Zhao (2022) Things not Written in Text: Exploring Spatial Commonsense from Visual Signals. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pp. 2365–2376. Cited by: §2.
  • F. Meng, W. Shao, L. Luo, Y. Wang, Y. Chen, Q. Lu, Y. Yang, T. Yang, K. Zhang, Y. Qiao, et al. (2024) PhyBench: A Physical Commonsense Benchmark for Evaluating Text-to-Image Models. CoRR. Cited by: §2.
  • L. Mitxelena (1981) Lengua común y dialectos vascos. Anuario del Seminario de Filología Vasca" Julio de Urquijo" 15, pp. 289–313. Cited by: §1.
  • S. Motamed, L. Culp, K. Swersky, P. Jaini, and R. Geirhos (2025) Do Generative Video Models Understand Physical Principles?. arXiv preprint arXiv:2501.09038. Cited by: §2.
  • R. Orlando, L. Moroni, P. Huguet Cabot, S. Conia, E. Barba, S. Orlandini, G. Fiameni, and R. Navigli (2024) Minerva LLMs: the first family of large language models trained from scratch on Italian data. In Proceedings of the Tenth Italian Conference on Computational Linguistics (CLiC-it 2024), F. Dell’Orletta, A. Lenci, S. Montemagni, and R. Sprugnoli (Eds.), pp. 707–719. Cited by: §4.3.
  • E. Pan, A. S. G. Choi, M. ter Hoeve, S. Seto, and A. Koenecke (2025) Analyzing Dialectical Biases in LLMs for Knowledge and Reasoning Benchmarks. arXiv preprint arXiv:2510.00962. Cited by: §2.
  • G. Pensa, B. Altuna, and I. Gonzalez-Dios (2024a) A Multi-layered Approach to Physical Commonsense Understanding: Creation and Evaluation of an Italian Dataset. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pp. 819–831. Cited by: §1, §1, §2, §2, §3.1, §3, §4.3.
  • G. Pensa, E. Azurmendi, J. Etxaniz, B. Altuna, and I. Gonzalez-Dios (2024b) GITA4CALAMITA-Evaluating the Physical Commonsense Understanding of Italian LLMs in a Multi-layered Approach: A CALAMITA Challenge. In Proceedings of the 10th Italian Conference on Computational Linguistics (CLiC-it 2024), pp. 1153–1160. Cited by: §1, §4.1.1, §4.1, §4.2, §4.3, §6.
  • M. Polignano, P. Basile, and G. Semeraro (2024) Advanced natural-based interaction for the italian language: llamantino-3-anita. ArXiv abs/2405.07101. Cited by: §4.3.
  • N. F. Rajani, B. McCann, C. Xiong, and R. Socher (2019) Explain Yourself! Leveraging Language Models for Commonsense Reasoning. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, A. Korhonen, D. Traum, and L. Màrquez (Eds.), Florence, Italy, pp. 4932–4942. Cited by: §2.
  • N. F. Rajani, R. Zhang, Y. C. Tan, S. Zheng, J. Weiss, A. Vyas, A. Gupta, C. Xiong, R. Socher, and D. Radev (2020) ESPRIT: Explaining Solutions to Physical Reasoning Tasks. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, D. Jurafsky, J. Chai, N. Schluter, and J. Tetreault (Eds.), Online, pp. 7906–7917. Cited by: §2.
  • O. Sainz, N. Pérez, J. Etxaniz, J. F. de Landa, I. Aldabe, I. García-Ferrero, A. Zabala, E. Azurmendi, G. Rigau, E. Agirre, M. Artetxe, and A. Soroa (2025) Instructing large language models for low-resource languages: a systematic study for basque. ArXiv abs/2506.07597. Cited by: §3.3, §4.3.
  • S. Storks, Q. Gao, and J. Y. Chai (2019) Commonsense Reasoning for Natural Language Understanding: A Survey of Benchmarks, Resources, and Approaches. arXiv preprint arXiv:1904.01172, pp. 1–60. Cited by: §2, §3, §4.3.
  • S. Storks, Q. Gao, Y. Zhang, and J. Chai (2021) Tiered Reasoning for Intuitive Physics: Toward Verifiable Commonsense Language Understanding. In Findings of the Association for Computational Linguistics: EMNLP 2021, pp. 4902–4918. Cited by: §1, §2, §4.2.
  • J. Sun, C. Zheng, E. Xie, Z. Liu, R. Chu, J. Qiu, J. Xu, M. Ding, H. Li, M. Geng, et al. (2025) A Survey of Reasoning with Foundation Models: Concepts, Methodologies, and Outlook. ACM Computing Surveys 57 (11), pp. 1–43. Cited by: §1.
  • G. Team, M. Riviere, S. Pathak, P. G. Sessa, C. Hardin, S. Bhupatiraju, L. Hussenot, T. Mesnard, B. Shahriari, A. Ramé, et al. (2024) Gemma 2: improving open language models at a practical size. arXiv preprint arXiv:2408.00118. Cited by: §4.3.
  • L. Uria and R. Etxepare (2012) Hizkeren arteko aldakortasun sintaktikoa aztertzeko metodologiaren nondik norakoak: basyque aplikazioa. Lapurdum. Euskal ikerketen aldizkaria| Revue d’études basques| Revista de estudios vascos| Basque studies review (16), pp. 117–135. Cited by: §2.
  • Y. Wang, J. Duan, D. Fox, and S. Srinivasa (2023) NEWTON: Are Large Language Models Capable of Physical Reasoning?. In Findings of the Association for Computational Linguistics: EMNLP 2023, H. Bouamor, J. Pino, and K. Bali (Eds.), Singapore, pp. 9743–9758. External Links: Link, Document Cited by: §2.
  • L. Weihs, A. Yuile, R. Baillargeon, C. Fisher, G. Marcus, R. Mottaghi, and A. Kembhavi (2022) Benchmarking Progress to Infant-level Physical Reasoning in AI. Transactions on Machine Learning Research. Cited by: §2.
  • S. Yu, P. Wu, P. P. Liang, R. Salakhutdinov, and L. Morency (2022) PACS: A Dataset for Physical Audiovisual CommonSense Reasoning. In European Conference on Computer Vision, pp. 292–309. Cited by: §2.
  • C. Ziems, W. Held, J. Yang, J. Dhamala, R. Gupta, and D. Yang (2023) Multi-VALUE: A Framework for Cross-Dialectal English NLP. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), A. Rogers, J. Boyd-Graber, and N. Okazaki (Eds.), Toronto, Canada, pp. 744–768. External Links: Link, Document Cited by: §2.
  • K. Zuazu (2008) Euskalkiak. euskararen dialektoak. Elkar. External Links: ISBN 978-84-90272-38-1 Cited by: §2.
Appendix A Automatic Adaptation Prompt
I will give you three versions of a story. Each version has five sentences. Some sentences are identical across versions. You need to adapt this text so that it includes Bizkaian dialectal features. You can use non-standard orthography. Try to make it as similar as possible to oral language.
Task:
  1. 1.
    First, list all unique sentences across all three stories.
  2. 2.
    Adapt each unique sentence exactly once into the Bizkaian dialect.
  3. 3.
    Then reconstruct the three stories with the translations, making sure that any identical source sentence always has the identical translation.
  4. 4.
    If there are more than three stories, repeat the same process for all of them.
Format:
This is an example of an standard (INPUT) instance and an example of the dialectal (OUTPUT) adaptation that you need to do:
Standard:
STORY1: [’Mikel lanera joan da’, ’Mikelek ordenagailua piztu du’, ’Mikelek mezuak irakurri ditu’, ’Mikelek mezuak erantzun ditu’, ’Mikel etxera joan da’]
STORY2: [’Mikel lanera joan da’, ’Mikelek mezuak erantzun ditu’, ’Mikelek mezuak irakurri ditu’, ’Mikelek ordenagailua piztu du’, ’Mikel etxera joan da’]
STORY3: [’Mikel lanera joan da’, ’Mikelek ordenagailua itzali du’, ’Mikelek mezuak irakurri ditu’, ’Mikelek mezuak erantzun ditu’, ’Mikel etxera joan da’]
Dialectal:
STORY1: [’Mikel lanera jun de’, ’Mikelek ordenagaillua piztu dau’, ’Mikelek mesuek irakurri dauz’, ’Mikelek mesuek erantzun dauz’, ’Mikel etxera jun de’]
STORY2: [’Mikel lanera jun de’, ’Mikelek mesuek erantzun ditu’, ’Mikelek mesuek irakurri dauz’, ’Mikelek ordenagaillua piztu dau’, ’Mikel etxera jun de’]
STORY3: [’Mikel lanera jun de’, ’Mikelek ordenagaillua amatatu dau’, ’Mikelek mesuek irakurri dauz’, ’Mikelek mesuek erantzun dauz’, ’Mikel etxera jun de’]
Output only the reconstructed stories in the exact same format as the input. Do not output explanations, steps, or commentary.